a boat heading due north crosses a wide river with speed of 12km/h relative to the water the water in the river has a uniform speed of 5 km/h due east relative to the earth.determine the velocity of the boat relative to an observer standing on either bank and the direction of boat
Answer #1
Do vector addition of the two velocities to obtain boat's velocity V w.r.t.river bank:

V = sqrt(12^2+5^2) = 13 Km/hr
and
tan theta = 5/12; (theta - angle between V and north)
answered by: tiffany jasper
Answer #2

Use Pythagorean theorem:

image.png


The boat relative to the observer is going at 13 km/h




answered by: 0bliviant